During today's maintenance we complete the installation of the NCal system in the North End building. More specifically:
- We brought R4-06 and R4-31 to the site and installed them respectively on the East Middle slot and the South Near slot.
- We moved R4-04 to the East Near slot (see picture East_setup), R4-05 to the South Middle slot, R4-02 to the North Far slot.
- We performed the zeroing of the three setups using new 2.5 mm calibrated spacers (see first picture). The residual offset was below 0.01 mm.
- We replaced the 2.5 mm spacers by 1 mm spacers, inducing a 1.5 mm offset, and observed the following offsets on the axial position: 1.56 mm for NSF, 1.53 mm for NSN, 1.54 mm for NEF, 1.47 mm for NEN, 1.58 mm for NNF and 1.52 mm for NNN.
- The setups have then been released and we roughly balanced and aligned them, the three figures (North, East and South) are showing the typical motion around their rest position.
- The cabling and configuration of the DAQ box 119 has been updated to include the NCal box temperature readout. For this purpose the two TCS readout have been moved to the Far end of this DAQ box (see photo of the new layout).
We left the power supplies on for the three setups.